# Synthesising Counterfactual Explanations via Label-Conditional Gaussian Mixture Variational Autoencoders

### v1

@@ 1 Introduction
In addition to these standard properties, various forms of robustness have been identified as a critical challenge (Upadhyay et al., 2021; Dominguez-Olmedo et al., 2022; Jiang et al., 2023; 2024a; Leofante & Wicker, 2025). Research in this area has been fragmented, often addressing specific robustness types in isolation or without considering their interplay with other key criteria (Jiang et al., 2024b). This has resulted in a lack of methods that can generate CEs integrating multiple properties, such as validity, plausibility, diversity, and multiple forms of robustness, within a single framework. This paper introduces a novel approach to bridge this gap, with a specific focus on (i) encouraging robustness to model changes, such that the recourse stays valid under model parameter changes (such as regular retraining), and (ii) guaranteeing robustness to input perturbations, which amounts to the stability of generated CEs against small perturbations applied to the input (Artelt et al., 2021; Leofante & Potyka, 2024). Indeed, if a method finds CEs that are readily invalidated under some model updates, or provides significantly different CEs for nearly identical inputs, the CEs are likely based on arbitrary model behaviour and algorithmic artefacts, rather than meaningful features. This particularly undermines CEs’ trustworthiness for recourse purposes.
A recent state-of-the-art method enhances robustness to input changes by leveraging diverse instances of nearest neighbours in the training dataset from the target class (Leofante & Potyka, 2024). However, this method remains heuristic and cannot guarantee perfect stability given the randomness in its distance-based thresholding process, and risks exposing sensitive training data. We argue that a more principled approach to robustness involves identifying a set of diverse, prototypical recourse points for the target class and then guiding all generated CEs to converge towards these points. While prior works have used generative models like standard variational autoencoders (VAEs), normalising flows, or diffusion models with similar intuitions, their primary goal was to improve plausibility by obtaining a realistic data manifold (Pawelczyk et al., 2020a; Wielopolski et al., 2024; Na & Lee, 2025). A key limitation is that they are typically unconditional, ignoring the information encoded in a classifier’s predicted labels that are readily available during recourse generation. As a result, sampling with randomness, or complex algorithms such as latent space gradient optimisation, is needed to perform the CE search in the latent space, without guaranteeing their validity.
@@ 2 Related Work
Counterfactual explanations are typically computed using specialised algorithms with gradient descent (Wachter et al., 2017) or mixed integer programming (Mohammadi et al., 2021) to address validity and proximity. To enhance these explanations, plausibility is often enforced by aligning CEs with the data manifold through nearest-neighbour approaches (Brughmans et al., 2023; Poyiadzi et al., 2020) or by using generative models (Mahajan et al., 2019; Pawelczyk et al., 2020a; Wielopolski et al., 2024; Na & Lee, 2025). Diversity can be incorporated through multi-objective optimisation (Mothilal et al., 2020; Dandl et al., 2020) or creating variations in a certain sampling process (Leofante & Potyka, 2024). More recently, various forms of robustness have been identified as a critical property of CEs (Mishra et al., 2021; Jiang et al., 2024b). These include robustness to input changes, ensuring similar inputs receive similar explanations (Artelt et al., 2021), and robustness to model changes, which requires CEs to remain valid after model updates like retraining or parameter shifts (Upadhyay et al., 2021; Dutta et al., 2022; Jiang et al., 2024c). This is often encouraged by pushing the CE further away from the classifier’s decision boundary into some more plausible regions (Pawelczyk et al., 2020b). We refer the reader to recent surveys (Guidotti, 2022; Karimi et al., 2023; Laugel et al., 2023; Jiang et al., 2024b) for in-depth reviews. Unlike most existing methods which address these properties in isolation, our proposed method is designed to simultaneously address a comprehensive set of these properties within a single, unified framework.
@@
Learning latent space with clusters. Early approaches, such as (Xie et al., 2016), learn low-dimensional embeddings for clustering using an autoencoder structure but lacked a generative component for sampling new data. VAEs (Kingma & Welling, 2014) provide a natural framework for this task by imposing a Gaussian prior on the latent space variables. The M2 model by Kingma et al. (2014) extends this idea to semi-supervised learning, formulating a mixture VAE where each latent component is explicitly tied to a class label. Building on this, the Gaussian Mixture VAE (GMVAE) (Dilokthanakul et al., 2016) introduces a categorical latent variable to model cluster assignments, thus encouraging structured latent spaces. A simplified variant by Shu (2016) incorporates the assignment variable directly into the encoder, yielding more effective clustering. In parallel, Conditional VAEs (Sohn et al., 2015) generalise the framework by conditioning both inference and generation on auxiliary attributes. Our proposed L-GMVAE combines these perspectives: it learns a Gaussian mixture–structured latent space while conditioning on predicted label information.
@@ 3.1 Label-Conditional GMVAE
L-GMVAE. Throughout the CE generation pipeline, the classifier-predicted label for the input and the desirable label for CEs are known. For computing CEs, it would therefore make sense to explicitly incorporate label information into the cluster assignments. Following GMVAE ((Shu, 2016)’s version, see Appendix A), the latent variable $z\in\mathbb{R}^{h}$ governs the data generation process. $z$ is regulated by a Gaussian Mixture with $K$ components, with a discrete variable $c\in\mathcal{C}=\{1,\ldots,K\}$ controlling the cluster assignment. In L-GMVAE, we partition the Gaussian clusters $\mathcal{C}$ using the possible $L$ labels in $\mathcal{Y}$ as: $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup\ldots\cup\mathcal{C}_{L}$, where each $\mathcal{C}_{i}\subset\mathcal{C}$, and for all $\mathcal{C}_{i},\mathcal{C}_{j}$ with $i\neq j$ and $i,j\in\mathcal{Y}$, $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$. The most intuitive approach for partitioning is to uniformly assign $L/K$ clusters for each class. Then, for a given class label $y$, its corresponding clusters are $\mathcal{C}_{y}$.
@@
$p(x,c,z\mid y)$ $=p(c\mid y)\,p(z\mid c)\,p(x\mid z)$ (1a)
$p(c\mid y)$ $=\frac{1}{|\mathcal{C}_{y}|}\text{ if }c\in\mathcal{C}_{y}$ (1b)
$p(z\mid c)$ $=\mathcal{N}(z\mid\mu_{z}(c),\sigma_{z}^{2}(c))$ (1c)
$p(x\mid z)$ $=\mathcal{N}(x\mid\mu_{x}(z),\sigma_{x}^{2}(z))$ (1d)
@@
$q(z,c\mid x,y)=q(c\mid x,y)\>q(z\mid x,c,y)$ (2)
We derive the evidence lower bound (ELBO) for this model, starting with the data log likelihood:
$log\>p(x)$ $=log\>\int_{z}\sum_{c\in\mathcal{C}_{y}}\>p(x,z,c\mid y)\>dz=log\>\int_{z}\sum_{c\in\mathcal{C}_{y}}\>q(z,c\mid x,y)\>\frac{p(x,z,c\mid y)}{q(z,c\mid x,y)}\>dz$ (3)
@@
$=\mathbb{E}_{q(z,c\mid x,y)}\left[\underbrace{\log\frac{p(c\mid y)}{q(c\mid x,y)}}_{\text{-KL($c$)}}+\underbrace{\log\frac{p(z\mid c)}{q(z\mid x,c,y)}}_{\text{-KL($z$)}}+\underbrace{\log\,p(x\mid z)}_{\text{Reconstruction}}\right]$
$\coloneq L_{\text{ELBO, LGMVAE}}$
Model architecture and training. In practice, the L-GMVAE consists of two neural network components, the inference model and the generative model, matching the above definitions. The inference model takes $x$ and the known $y$ label as inputs, and outputs the probability distributions over the clusters. These are then concatenated together to predict $z$. The generative model produces a reconstructed input using $z$, and learns the Gaussian mixture prior ($p(z\mid c)$). The ELBO objective has three terms, Kullback–Leibler (KL) divergence for $c$ and $z$ as regularisers, and a reconstruction term. The KL terms are expressed analytically, while the reconstruction term is the MSE loss between $x$ and the reconstructed $x^{\prime}$. The practical loss function is a negated, weighted sum of these ELBO terms. See Appendices B and C for more details of our trained L-GMVAE models.
@@
This characteristic motivates using these reconstructed centroids as part of the CE generation method, because they guarantee the validity of the generated CEs. Furthermore, these points should be well within the data manifold. Therefore, if a classifier is updated, these prototypes are more likely to retain their predicted class, promoting plausibility and robustness to model changes for CEs. Finally, if we enforce that all CEs converge to the vicinity of the centroids of one class, it would also enhance the robustness of CEs to input changes. These properties are all quantitatively measured in our experiments in Section 4.
@@ 3.2 Latent Paths and Counterfactual Explanations
Next, we describe how L-GMVAE can be used to generate paths of counterfactual explanations leading to high-quality recourse. After learning, the L-GMVAE’s latent cluster centroids are fixed and available through the optimised prior distribution $p(z\mid c)$ by taking the mean of the Gaussian component, denoted as $z_{c_{i}}$ for each $c_{i}\in\mathcal{C}$. We refer to the L-GMVAE’s decoder reconstruction functionality as $Dec:\mathbb{R}^{h}\rightarrow\mathbb{R}^{d}$. Then, for a target counterfactual class $y^{\prime}$ and its associated clusters $\mathcal{C}_{y^{\prime}}$, $Dec(z_{c_{j}})$ is the reconstructed cluster centroid for each $c_{j}\in\mathcal{C}_{y^{\prime}}$. From the discussions in Section 3.1, these $Dec(z_{c_{j}})$ points carry suitable properties for recourse purposes because they are valid, plausible, robust, and diverse. Additionally, they do not risk exposing existing data points given their synthetic nature.
@@ 3.3 Accommodating Actionability Constraints
LAPACE also supports actionability constraints. We define a set of constraints in the input space using a differentiable function $g(x)$, where $g(x)\leq 0$ indicates the constraint is satisfied. For example, for a feature $i$ that must not be greater than or equal to a value $A$, the function is simply $g(x)=x_{i}-A$. Constraints on multiple features can be intuitively summed together.
@@ 4.1 Experiment Setup
Datasets and classifiers. We run our experiments on four datasets, heloc credit FICO (2018), wine quality Cortez et al. (2009), adult income Becker & Kohavi (1996), and compas recidivism Julia Angwin & Kirchner (2016). We split each dataset into a train and a test set for training classifiers. To demonstrate the model-agnostic nature of our approach, we train a random forest and a neural network for each dataset, which we use to compute the CEs. Characteristics of the datasets and models are summarised in Table 1.
@@
CE evaluation procedure and metrics [^1] The evaluation metrics and baseline method implementations are adapted from CARLA library (Pawelczyk et al., 2021) and RobustX library (Jiang et al., 2025). All experiments were performed on a linux machine with an Intel Xeon w5-2455X CPU with 128GB RAM, and 2x24GB NVIDIA GeForce RTX 4090 GPUs.. For each dataset-classifier pair, we obtain a CE test set of 100 test points predicted with label 0. We then use each method to generate CEs and evaluate on the following metrics. This process is repeated 5 times on different test sets. The procedures are in line with the existing literature. Validity: the portion of test points receiving valid CEs. Proximity: the average L1 distance between a test point and its CE(s). Lower distances indicate that the recourse prescribed by the CE would be easier to achieve by the end user, therefore it is more preferred (Wachter et al., 2017). Plausibility: the average local outlier factor (lof) (Breunig et al., 2000) of each CE. Lof identifies outliers by measuring the local density deviation of a data point with respect to its neighbours, and lower values indicate better proximities to realistic data, therefore better plausibility. Diversity: if a CE algorithm finds multiple CEs per test input, diversity is calculated as the average pair-wise (L1) distance between these CEs. Higher distances indicate that the CEs are more diverse. Robustness to Model Changes: we obtain 20 retrained classifiers using different selections of the dataset as the training set, and test on average how many CEs are valid on these new classifiers. Robustness to Input Changes evaluates the stability of generated CEs when small norm-bounded perturbations are applied to the test input. We perturb the test input 10 times with small noises, generate new CEs for each of the perturbations, and report the average distance between the new CEs and the original CE (or max-set-distance (Leofante & Potyka, 2024) if the method supports finding multiple CEs per input). Smaller distances would mean better stability - finding similar CEs for similar inputs.
Additionally, on the two continuous-only datasets, we manually specify some actionability constraints (10 and 18 for wine and heloc) in the style of previous work in tabular data generation (Stoian & Giunchiglia, 2025). Each constraint specifies either a range for a feature value or the relation between two features (one greater than another) for the resulting CE. We randomly select 5 constraints for each test input, which are then enforced on a path of CEs and check if there is at least one CE satisfying all constraints while staying valid.
@@
DRCE (Leofante & Potyka, 2024) -

### v2

@@ 1 Introduction
In addition to these standard properties, various forms of robustness have been identified as a critical challenge (Upadhyay et al., 2021; Dominguez-Olmedo et al., 2022; Jiang et al., 2023; 2024b; Leofante and Wicker, 2025). Research in this area has been fragmented, often addressing specific robustness types in isolation or without considering their interplay with other key criteria (Jiang et al., 2024c). This has resulted in a lack of methods that can generate CEs integrating multiple properties, such as validity, plausibility, diversity, and multiple forms of robustness, within a single framework. This paper introduces a novel approach to bridge this gap, with a specific focus on (i) encouraging robustness to model changes, such that the recourse stays valid under model parameter changes (such as regular retraining), and (ii) guaranteeing robustness to input perturbations, which amounts to the stability of generated CEs against small perturbations applied to the input (Artelt et al., 2021; Leofante and Potyka, 2024). Indeed, if a method finds CEs that are readily invalidated under some model updates, or provides significantly different CEs for nearly identical inputs, the CEs are likely based on arbitrary model behaviour and algorithmic artefacts, rather than meaningful features. This particularly undermines CEs’ trustworthiness for recourse purposes.
A recent state-of-the-art method enhances robustness to input changes by leveraging diverse instances of nearest neighbours in the training dataset from the target class (Leofante and Potyka, 2024). However, this method remains heuristic and cannot guarantee perfect stability given the randomness in its distance-based thresholding process, and risks exposing sensitive training data. We argue that a more principled approach to robustness involves identifying a set of diverse, prototypical recourse points for the target class and then guiding all generated CEs to converge towards these points. While prior works have used generative models like standard variational autoencoders (VAEs), normalising flows, or diffusion models with similar intuitions, their primary goal was to improve plausibility by obtaining a realistic data manifold (Pawelczyk et al., 2020a; Wielopolski et al., 2024; Na and Lee, 2025). A key limitation is that they are typically unconditional, ignoring the information encoded in a classifier’s predicted labels that are readily available during recourse generation. As a result, sampling with randomness, or complex algorithms such as latent space gradient optimisation, is needed to perform the CE search in the latent space, without guaranteeing their validity.
@@ 2 Related Work
Counterfactual explanations are typically computed using specialised algorithms with gradient descent (Wachter et al., 2017) or mixed integer programming (Mohammadi et al., 2021) to address validity and proximity. To enhance these explanations, plausibility is often enforced by aligning CEs with the data manifold through nearest-neighbour approaches (Brughmans et al., 2023; Poyiadzi et al., 2020) or by using generative models (Mahajan et al., 2019; Pawelczyk et al., 2020a; Wielopolski et al., 2024; Na and Lee, 2025). Diversity can be incorporated through multi-objective optimisation (Mothilal et al., 2020; Dandl et al., 2020) or creating variations in a certain sampling process (Leofante and Potyka, 2024). More recently, various forms of robustness have been identified as a critical property of CEs (Mishra et al., 2021; Jiang et al., 2024c). These include robustness to input changes, ensuring similar inputs receive similar explanations (Artelt et al., 2021), and robustness to model changes, which requires CEs to remain valid after model updates like retraining or parameter shifts (Upadhyay et al., 2021; Dutta et al., 2022; Jiang et al., 2024a). This is often encouraged by pushing the CE further away from the classifier’s decision boundary into some more plausible regions (Pawelczyk et al., 2020b). We refer the reader to recent surveys (Guidotti, 2022; Karimi et al., 2023; Laugel et al., 2023; Jiang et al., 2024c) for in-depth reviews. Unlike most existing methods which address these properties in isolation, our proposed method is designed to simultaneously address a comprehensive set of these properties within a single, unified framework.
Parallel to the CE research in the explainable AI literature which concerns revealing faithfully what changes would be needed for a given classifier to flip its prediction, there have been substantial research efforts on causal inference, that instead aim at causally characterising the underlying data generation process (Pawlowski et al., 2020; Shen et al., 2022; Ribeiro et al., 2023; Monteiro et al., 2023; Wu et al., 2025; Ribeiro et al., 2025). A causal model is often explicitly considered, allowing counterfactual inference (Pearl, 2009) and thereby discovering causes behind certain observations (Halpern, 2016). From a cognitive science perspective, CEs have been shown to serve complementary goals for end-user perception compared to the causal explanations (Byrne, 2019), highlighting the inherent differences between the two fields. Our work follows the standard CE setups (causality-agnostic) introduced above, and incorporating causal constraints into the CE generation process (as studied by Karimi et al. (2020; 2021)) would be exciting future work.
@@
Learning latent space with clusters. Early approaches, such as (Xie et al., 2016), learn low-dimensional embeddings for clustering using an autoencoder structure but lacked a generative component for sampling new data. VAEs (Kingma and Welling, 2014) provide a natural framework for this task by imposing a Gaussian prior on the latent space variables. The M2 model by Kingma et al. (2014) extends this idea to semi-supervised learning, formulating a mixture VAE where each latent component is explicitly tied to a class label. Building on this, the Gaussian Mixture VAE (GMVAE) (Dilokthanakul et al., 2016) introduces a categorical latent variable to model cluster assignments, thus encouraging structured latent spaces. A simplified variant by Shu (2016) incorporates the assignment variable directly into the encoder, yielding more effective clustering. In parallel, Conditional VAEs (Sohn et al., 2015) generalise the framework by conditioning both inference and generation on auxiliary attributes. Our proposed L-GMVAE combines these perspectives: it learns a Gaussian mixture–structured latent space while conditioning on predicted label information.
@@ 3.1 Label-Conditional GMVAE
L-GMVAE. Throughout the CE generation pipeline, the classifier-predicted label for the input and the desirable label for CEs are known. For computing CEs, it would therefore make sense to explicitly incorporate label information into the cluster assignments. Following GMVAE ((Shu, 2016)’s version, see Appendix A), the latent variable $z\in\mathbb{R}^{h}$ governs the data generation process. $z$ is regulated by a Gaussian Mixture with $K$ components, with a discrete variable $c\in\mathcal{C}=\{1,\ldots,K\}$ controlling the cluster assignment. In L-GMVAE, we partition the Gaussian clusters $\mathcal{C}$ using the possible $L$ labels in $\mathcal{Y}$ as: $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup\ldots\cup\mathcal{C}_{L}$, where each $\mathcal{C}_{i}\subset\mathcal{C}$, and for all $\mathcal{C}_{i},\mathcal{C}_{j}$ with $i\neq j$ and $i,j\in\mathcal{Y}$, $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$. The most intuitive approach for partitioning is to uniformly assign $K/L$ clusters for each class. Then, for a given class label $y$, its corresponding clusters are $\mathcal{C}_{y}$.
@@
$p(x,c,z\mid y)$ $=p(c\mid y)\,p_{\theta}(z\mid c)\,p_{\theta}(x\mid z)$ (1a)
$p(c\mid y)$ $=1/{|\mathcal{C}_{y}|}\text{ if }c\in\mathcal{C}_{y}$ (1b)
$p_{\theta}(z\mid c)$ $=\mathcal{N}(z\mid\mu_{z}(c),\sigma_{z}^{2}(c))$ (1c)
$p_{\theta}(x\mid z)$ $=\mathcal{N}(x\mid\mu_{x}(z),\sigma_{x}^{2}(z))$ (1d)
@@
$q(z,c\mid x,y)=q_{\phi}(c\mid x,y)\>q_{\phi}(z\mid x,c,y)$ (2)
where $p_{\theta}$ and $q_{\phi}$ indicate that relevant components are approximated via neural networks. We derive the evidence lower bound (ELBO) for this model, starting with the data log likelihood:
$log\>p(x|y)$ $=log\>\int_{z}\sum_{c\in\mathcal{C}_{y}}\>p(x,z,c\mid y)\>dz=log\>\int_{z}\sum_{c\in\mathcal{C}_{y}}\>q(z,c\mid x,y)\>\frac{p(x,z,c\mid y)}{q(z,c\mid x,y)}\>dz$ (3)
@@
$=\mathbb{E}_{q(z,c\mid x,y)}\left[\underbrace{\log\frac{p(c\mid y)}{q_{\phi}(c\mid x,y)}}_{\text{-KL($c$)}}+\underbrace{\log\frac{p_{\theta}(z\mid c)}{q(z_{\phi}\mid x,c,y)}}_{\text{-KL($z$)}}+\underbrace{\log\,p_{\theta}(x\mid z)}_{\text{Reconstruction}}\right]$
$\coloneq L_{\text{ELBO, L-GMVAE}}$
Model architecture and training. In practice, the L-GMVAE consists of two neural network components, the inference model $q_{\phi}$ and the generative model $p_{\theta}$, matching the above definitions. The inference model takes $x$ and the known $y$ label as inputs, and outputs the probability distributions over the clusters. These are then concatenated together to predict $z$. The generative model produces a reconstructed input using $z$, and learns the Gaussian mixture prior ($p_{\theta}(z\mid c)$). The ELBO objective has three terms, Kullback–Leibler (KL) divergence for $c$ and $z$ as regularisers, and a reconstruction term. The KL terms are expressed analytically, while the reconstruction term is the MSE loss between $x$ and the reconstructed $x^{\prime}$. The practical loss function is a negated, weighted sum of these ELBO terms. See Appendices B and C for more details of our trained L-GMVAE models.
@@
This characteristic motivates using these reconstructed centroids as part of the CE generation method, because they guarantee the validity of the generated CEs. Furthermore, these points should be well within the data manifold. Therefore, if a classifier is updated, these prototypes are more likely to retain their predicted class, promoting plausibility and robustness to model changes for CEs (Pawelczyk et al., 2020b). Finally, if we enforce that all CEs converge to the vicinity of the centroids of one class, it would also enhance the robustness of CEs to input changes. These properties are all quantitatively measured in our experiments in Section 4.
@@ 3.2 Latent Paths and Counterfactual Explanations
Next, we describe how L-GMVAE can be used to generate paths of counterfactual explanations leading to high-quality recourse. After learning, the L-GMVAE’s latent cluster centroids are fixed and available through the optimised prior distribution $p_{\theta}(z\mid c)$ by taking the mean of the Gaussian component, denoted as $z_{c_{i}}$ for each $c_{i}\in\mathcal{C}$. We refer to the L-GMVAE’s decoder reconstruction functionality as $Dec:\mathbb{R}^{h}\rightarrow\mathbb{R}^{d}$. Then, for a target counterfactual class $y^{\prime}$ and its associated clusters $\mathcal{C}_{y^{\prime}}$, $Dec(z_{c_{j}})$ is the reconstructed cluster centroid for each $c_{j}\in\mathcal{C}_{y^{\prime}}$. From the discussions in Section 3.1, these $Dec(z_{c_{j}})$ points carry suitable properties for recourse purposes because they are valid, plausible, robust, and diverse. Additionally, they do not risk exposing existing data points given their synthetic nature.
@@ 3.3 Accommodating Actionability Constraints
[…]
@@ 4.1 Experiment Setup
Datasets and classifiers. We run our experiments on datasets which are commonly used in prior work (Karimi et al., 2023) and are advocated in the CE benchmarking library, CARLA (Pawelczyk et al., 2021). They are heloc credit (FICO, 2018), wine quality (Cortez et al., 2009), adult income (Becker and Kohavi, 1996), and compas recidivism (Julia Angwin and Kirchner, 2016). We split each dataset into a train and a test set for training classifiers. To demonstrate the model-agnostic nature of our approach, we train a random forest and a neural network for each dataset, which we use to compute the CEs. Characteristics of the datasets and models are summarised in Table 1.
@@
CE evaluation procedure and metrics [^1] The evaluation metrics and baseline method implementations are adapted from CARLA library (Pawelczyk et al., 2021) and RobustX library (Jiang et al., 2025). All experiments were performed on a linux machine with an Intel Xeon w5-2455X CPU with 128GB RAM, and 2x24GB NVIDIA GeForce RTX 4090 GPUs.. For each dataset-classifier pair, we obtain a CE test set of 100 test points predicted with label 0. We then use each method to generate CEs and evaluate on the following metrics. This process is repeated 5 times on different test sets. The procedures are in line with the existing literature. Validity: the portion of test points receiving valid CEs. Proximity: the average L1 distance between a test point and its CE(s). Lower distances indicate that the recourse prescribed by the CE would be easier to achieve by the end user, therefore it is more preferred (Wachter et al., 2017). Plausibility: the average local outlier factor (lof) (Breunig et al., 2000) of each CE. Lof identifies outliers by measuring the local density deviation of a data point with respect to its neighbours, and lower values indicate better proximities to realistic data, therefore better plausibility. Diversity: if a CE algorithm finds multiple CEs per test input, diversity is calculated as the average pair-wise (L1) distance between these CEs. Higher distances indicate that the CEs are more diverse. Robustness to Model Changes: we obtain 20 retrained classifiers using different selections of the dataset as the training set, and test on average how many CEs are valid on these new classifiers. Robustness to Input Changes evaluates the stability of generated CEs when small norm-bounded perturbations are applied to the test input. We perturb the test input 10 times with small noises, generate new CEs for each of the perturbations, and report the average distance between the new CEs and the original CE (or max-set-distance (Leofante and Potyka, 2024) if the method supports finding multiple CEs per input). Smaller distances would mean better stability - finding similar CEs for similar inputs.
Additionally, on the two continuous-only datasets, we manually specify some actionability constraints (10 and 18 for wine and heloc) in the style of previous work in tabular data generation (Stoian and Giunchiglia, 2025). Each constraint specifies either a range for a feature value or the relation between two features (one greater than another) for the resulting CE. We randomly select 5 constraints for each test input, which are then enforced on a path of CEs and check if there is at least one CE satisfying all constraints while staying valid.
@@
DRCE (Leofante and Potyka, 2024) -